Write an equation in point-slope form for the line through the given points or through the given point with the given slope. (-2,3); m = 4
Answer by butterfliiizzz(18)

Given information:
(-2,3)
m=4
The point slope equation is y - y1 = m (x - x1)
The 1s in this equation are subscripts, meaning that they appear just below the variable, as opposed to an exponent, which is a superscript, which appears just above the variable.
So the point (-2,3) will be filled in where x1 and y1 occur in the equation. m is the slope.
So the equation will look like this:
y - 3 = 4 (x - -2)
minus a negative number will turn into a plus:
y - 3 = 4 (x + 2)
Now we will distribute the 4 by multiplying:
y - 3 = 4x + 8
We want to get the equation in y= mx +b format, so we need to get y by itself, so we will subtract 3 from each side (what you do to one side, you must do to the other):
y = 4x + 5